Systems Staff Meeting Notes 3/12/01

Systems Staff Meeting Notes
March 12, 2001

Present: Rich Bennett, Cynthia Bowker, Will Chaney, Bill Covey, Debra Harris, Winston Harris, Suzy Shaw, Hilton Soto, Geoffrey Stewart, Ying Tang
Announcements
  1. NT Fault Scan - look in the Systems folder on NT2 for the NT Fault Scan subfolder - we need to run the unzipped Patchwrk on any work station or server that is running NT. Just copy the unzipped version to the desktop, run it - then run the file scan.
  2. Cracks in the building - Rob Roberts said he would make the call urgent; if we haven't heard from anyone by Wednesday then we will call him again.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. SUSE Linux 7.1 Professional software has been received. Will is fast becoming an expert.
  2. 60 Dells received last week.
  3. Resource Services picked up 10 machines last week. Documents picked up quite a few as well.
Liaisons
  1. Setting up the new Dells - Geoffrey has noticed that the display prefers a refresh rate of 100 to be set or else it gives a wavy lined, strobing monitor effect. This is different than the jittery monitor problem that occurred in a prior shipment. All of these are set to a different default and will need to be corrected. Liaisons need to be notified.
Networking
  1. Power Outage in MSL over the weekend - Will came in on Saturday and got everything back up, except the breaker for John Ashcraft's office - it was off & he couldn't find it. Michael Howell came in on Sunday morning and got the breaker switched back on. Everything was fine.
  2. Wiring abatement - there is still some wiring left in the mezzanine area. They are scheduled to remove the remaining wiring in Preservation tomorrow.
  3. Wiring project in Journalism - should be finished tomorrow. The false wall and wires have been installed, they should be switched over to the new wiring by tomorrow.
  4. We are in negotiation with Gary Cornwell to set up a Circulation point in room 100 Library East. Wants to use it on nights and weekends only to get the Dewey 900's circulated. Will probably move the guard into room 100 to secure the work station as well as monitoring the noise complaints that have been occurring in that room after hours. Time table looks like a couple of months.
  5. Erich is anxious about the rest of his Apache servers. The SID server will be the Apache rather than NT8. It can really load pictures fast - and hold the resolution. It is a magical piece of software developed by NASA. It really optimizes the loading of Web pages - we may want to use it for other sites. Go to http://www.lizardtech.com to check out samples of the software. It is a plug-in, but we will run it on the server.
Software Projects Status
  1. Resource Services' PAULA - Winston says they have a version that's ready - Bill says let's get it out to them. Ask Diana Hagan how she wants to let it out; she will ultimately be in charge of monitoring its use in the department. Winston reports that Cynthia is working on a few other options, but this one should go now.
  2. OPS rewrite - still ongoing.
  3. Ringling - the two parts pf the project that Ying has finished were reviewed by Erich last Wednesday - he is pleased. She has two more parts left to do. She needs to let them enter data and search items by key words. May be easier to do in Access than in Delphi.
  4. ILLIAD integration with the Accounting system - is ready to go. It can be put on the SQL server, so we won't have to buy a separate server. Barbara Oliver is in the process of cutting a check for it.
DP Coordinator
  1. Students are back, with the exception of Melissa who should be returning from China today and back at work tomorrow.
  2. Tape dispersion ongoing.
  3. Calendar updates continue - be sure to review the calendar for upcoming outages.


Back to the Systems Homepage

Last updated March 19, 2001
by Debra Harris
debharr@mail.uflib.ufl.edu